OpenStudy (anonymous):

A train traveled 72 miles in 1.5 hours before making an emergency stop. After the stop, it traveled 48 miles in 2 hours. A. Write the formula that relates distance to rate and time. B. Use the formula to find the train’s average speed. Show your work. Round to the nearest tenth.

OpenStudy (anonymous):

A. r=d/t

OpenStudy (anonymous):

B. to do this plug in the total distance and the total time traveled for d and t respectively

OpenStudy (anonymous):

r = (72+48)/(2+1.5) = ?